OpenStudy (anonymous):

Simplify completely 4 m cubed over 3 m squared n times 9 n squared over 30 m squared n squared

OpenStudy (anonymous):

@ositan

OpenStudy (anonymous):

@april115

OpenStudy (anonymous):

please I really need help

OpenStudy (anonymous):

@jigglypuff314

OpenStudy (anonymous):

I will give medals

jigglypuff314 (jigglypuff314):

umm... I can try... but I'm kinda going cross-eyed >,< could you try pressing the "Equation" button or the "Draw" button to make it more clear? Thanks :)

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

okay cool, thanks :) now you know how m over m = 1 and therefore cancels out when multiplying?

OpenStudy (anonymous):

yes

jigglypuff314 (jigglypuff314):

then \[\frac{ 4m ^{3} }{ 3m ^{2}n } * \frac{ 9n ^{2} }{ 30m ^{2}n ^{2} } \rightarrow \frac{ 4m ^{2}9n ^{2} }{ 3m ^{2}n30m ^{2}n ^{2} }\]

jigglypuff314 (jigglypuff314):

and\[\frac{ 4m ^{2}9n ^{2} }{ 3m ^{2}n30m ^{2}n ^{2} } \rightarrow \frac{ 4*9*m*m*n*n }{ 3*30*m*m*m*m*n*n*n }\]

jigglypuff314 (jigglypuff314):

m*m*m can be canceled out from the top and bottom n*n can be canceled out from the top and bottom as well then you'll be left with: \[\frac{ 36 }{ 90mn }\]then 36 over 90 can be simplified :)

OpenStudy (anonymous):

the choices are 1) 2n -- 3m 2) 2m -- 5n 3) 2 -- 3mn 4) 2 -- 5mn

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

36 and 90 can both be divided by 18 so 36 over 90 would equal 2 over 5 and mn is in the denominator :)

OpenStudy (anonymous):

thank you so much!!!

jigglypuff314 (jigglypuff314):

I hope you understand it at least :/

OpenStudy (anonymous):

I do....its just a little confusing at first

jigglypuff314 (jigglypuff314):

yeah, okay, practice makes perfect! :) good luck and have a nice day!

OpenStudy (anonymous):

thank you, you too